OpenStudy (anonymous):

Solve the equation 2/5 x − 3 = 13 for x. Please show your work and solution in the space provided.

OpenStudy (zrsmith):

First, add 3 to both sides: 2/5x-3+3 = 13+3 =2/5x = 16 Next multiply both sides of the equation by 5 to get rid of the fraction. 5* 2/5x = 5*16 2x = 80 Next divide both sides by 2 2/2x = 80/2 EQUALS: x =40 Then substitute x to make sure the answer is correct: 2/5*40-3 =? 13 2*8 - 3 =? 13 16-3 = 13 13=13 The answer checks!

OpenStudy (anonymous):

thank you

OpenStudy (zrsmith):

you are welcome.

OpenStudy (anonymous):

can you please help me with a another one

OpenStudy (zrsmith):

i can try

OpenStudy (zrsmith):

what is the problem?